OpenStudy (anonymous):

PLEASE HELP! The area of a rectangle is 52ft^2 , and the length of the rectangle is 5ft less than twice the width. Find the dimensions of the rectangle.

jimthompson5910 (jim_thompson5910):

Let W be the width of the rectangle. It's unknown for now. "the length of the rectangle is 5ft less than twice the width" tells us the length is L = 2W - 5

jimthompson5910 (jim_thompson5910):

so we have this rectangle 

jimthompson5910 (jim_thompson5910):

What is the area of any rectangle in general?

OpenStudy (anonymous):

A=1/2*b*h

jimthompson5910 (jim_thompson5910):

That's for a triangle

jimthompson5910 (jim_thompson5910):

For a rectangle is A = L*W

OpenStudy (anonymous):

hahah yeah my bad

jimthompson5910 (jim_thompson5910):

In this specific problem, this means A = L*W 52 = (2W-5)*W 52 = W*(2W-5) 52 = 2W^2 - 5W 0 = 2W^2 - 5W - 52 2W^2 - 5W - 52 = 0 Now use the quadratic formula to solve for W

OpenStudy (anonymous):

what quadratic formula?

jimthompson5910 (jim_thompson5910):

Have you learned about this formula before? \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

yes, but how do i know which numbers to plug them into the letters

jimthompson5910 (jim_thompson5910):

2W^2 - 5W - 52 is the same as 2x^2 - 5x - 52 after you replace W with x

jimthompson5910 (jim_thompson5910):

2x^2 - 5x - 52 is in the form ax^2 + bx + c a = 2 b = -5 c = -52

jimthompson5910 (jim_thompson5910):

\[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[\Large x=\frac{-(-5) \pm \sqrt{(-5)^2-4(2)(-52)}}{2(2)}\] I'll let you finish